3u-7 + 7u-6 = 11
6 3
The 3u-7 is over the 6 and the 7u-6 is over the 3. I tried it with a common denominator of 6. The 7 and 11 is confusing me.
Thank you!

Answer by fractalier(6550)

Multiply everything by 6 in order to clear all fractions...we get
3u - 7 + 2(7u - 6) = 66
3u - 7 + 14u - 12 = 66
17u - 19 = 66
17u = 85
u = 5